OpenStudy (anonymous):

x^5-12x^3+32x=0 Please help and show work

OpenStudy (anonymous):

x(x^4-12x^2+32)=0 x((4+-1x)(8+-1x))=0

OpenStudy (solomonzelman):

x^5-12x^3+32x=0 x(x^4-12x^2+32)=0 x^4-12x^2+32=0 let x^2=a a^2-12a+32=0 Do it from here.

OpenStudy (anonymous):

Can you continue from there?

OpenStudy (anonymous):

how @SolomonZelman get X^2=a?

OpenStudy (solomonzelman):

I didn't get it, I am substituting a for x^2 so that it would be easier to solve. it's a make up.

OpenStudy (anonymous):

so either x=0, 4+-1x=0, or 8+-1x=0. so x {0,4,8}

OpenStudy (solomonzelman):

you didn't show any work, you do realize that.

OpenStudy (anonymous):

i'm confused

OpenStudy (solomonzelman):

x^5-12x^3+32x=0 factor out of x, x(x^4-12x^2+32)=0 divide both sides by x, x^4-12x^2+32=0 let x^2=a (substituting "a" for "x^2" ) it becomes, a^2-12a+32=0 can you solve this quadratic?

OpenStudy (solomonzelman):

Or you can forget about a=x^2, just say x^4-12x^2+32=0 (x^2-4)(x^2-8)=0 x^2-4=0 or x^2-8=0 x^2=4 or x^2=8\ x= +- 2 or x= +- 2 sqrt(2)

OpenStudy (anonymous):

I subtract 32 from zero first right

OpenStudy (anonymous):

oh so there are two answers?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I can't start from a scratch.

OpenStudy (anonymous):

huh?